Calculate point elasticity of demand for demand function Q=10-2p for decrease in price from Rs 3 to Rs 2
1
Expert's answer
2014-03-26T11:22:03-0400
Calculate point elasticity of demand for demand function Q=10-2p for decrease in price from Rs 3 to Rs 2
Solution
P1=3; Q1=10-2*P1=10-2*3=10-6=4
P2=2; Q2=10-2*P2=10-2*2=10-4=6; decrease in price ΔP=P2-P1=2-3=-1; decrease in demand ΔQ=Q2-Q1=6-4=2
Elasticity of demand is calculated by E=(P1+P2)*ΔQ/((Q1+Q2)*ΔP)=(3+2)*2/((4+6)*(-1))=-1